OpenStudy (anonymous):

Find the sum of the first n terms of the series: 3 + 7 + 13 + 21 + 31 + …

OpenStudy (hba):

Do you know the formula ?

OpenStudy (anonymous):

What shall I say .. Erm, I do, but this one's a bit different.. the differences of the terms of this series are in AP .. :/

OpenStudy (hba):

Do you know that \[a _{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

obviously I do :P

OpenStudy (hba):

Tell me what is a1 and what is d ?

OpenStudy (anonymous):

This series isn't an AP , no? I can't just take the first term and the c.d .. right?

OpenStudy (hba):

my bad :(

OpenStudy (cwrw238):

right its not an AP - though the differences are an AP

OpenStudy (hba):

@cwrw238 How do we change it into an AP then ?

OpenStudy (hba):

I do not really remember :(

OpenStudy (anonymous):

@cwrw238 - Exactly! @hba - Nevermind :)

OpenStudy (cwrw238):

good question hba - i must admit i've forgotten the method - i'll have to revise!!

OpenStudy (anonymous):

I don't even know the method :/

OpenStudy (hba):

Same here :(

OpenStudy (cwrw238):

i did it in GCSE level (aged 15) - darn it i remember a square term is involved

OpenStudy (hba):

Look it up,lets do some revision :D

OpenStudy (cwrw238):

yea!! the nth term of this sequence is n^2 + n + 1 - i think

OpenStudy (hba):

n^2+n-1 i guess lol

OpenStudy (cwrw238):

n=1 1 + 1 + 1 = 3 2 2^2 + 2 + 1 = 7 3 3^2 + 3 + 1 = 13

OpenStudy (cwrw238):

yea - i guessed right!

OpenStudy (cwrw238):

now we need the sum

OpenStudy (hba):

ok good :)

OpenStudy (cwrw238):

SIGMA n = n/2 [ n + 1] SIGMA 1 = n formula for sum of n^2 - we'd have to google that - there might be a better way to do this - lol

OpenStudy (cwrw238):

google has it (1/6) n(n+1)(2n+1)

OpenStudy (cwrw238):

so finally - its sum of n terms = (1/6) n(n+1)(2n+1) + (1/2)( n + 1) + n which can be simplified but i'm too tired!!

OpenStudy (cwrw238):

method to find nth term when difference is in AP: seq. no. n 1 2 3 4 n^2 1 4 9 16 n^2 + n 2 6 12 20 and as the sequence is 3,7,13 nth term is n^2 + n + 1

OpenStudy (cwrw238):

method to find nth term when difference is in AP: seq. no. n 1 2 3 4 n^2 1 4 9 16 n^2 + n 2 6 12 20 and as the sequence is 3,7,13 nth term is n^2 + n + 1

OpenStudy (cwrw238):

- i remembered it at last